uctor Check Compute the amounts that will be accumulated. Round to the nearest dollar. Use the appropriate future value table that may be found by dicking on the Future Value link. Bank 2 Check My W 0- Icon Ke
Answer #1

Solution:

Bank 1:

Amount invested in Bank 1 = $2,000

interest rate = 5% simple interest

Period of investment = 5 years

Interest earned in 5 years = P*R*T = $2,000*5%*5 = $500

Amount accumulated after 5 years = $2,000 + $500 = $2,500

?Bank 2:

Amount invested in Bank 2 = $2,000

interest rate = 5% compounded anually

Period of investment = 5 years

Future value of investment = P (1+r)^t

= $2,000 (1+0.05)^5 = $2,553

Total amount accumulated from bank 1 and bank 2 = $2,500 + $2,553 = $5,053